A rotating step shaft is loaded as shown, where the forces FA and FB are constant at 600 lbf and 300 lbf, respectively, and tF = 600 lbf 6 in Problem 7-5 Tin FR = 300 lbf
Answer #1

In the given problem load is applied on the beam in two different planes i.e. X-Y plane and Z-X plane. So, we will analyse thDraw the SFD, TMD, and BMD as follows: 400 lbf 0 0 200 lbf SFD T 0 0 TMD 2400 lb-in 1200 lb-in 0 0 BMD Shear force calculatioTo =TA = 0 Tg = 1 Tc =T-T=0 Bending moment calculation. = Mo = 0 MA= R,x6=400x6 = 2400 lb-in M = (400x12)-(600x6) = 1200 lb-iFor beam to be in equilibrium, Mo = 0 (300x12)-18R, = 0 R = 200 lbf Fz=0 R; +R, -400 = 0 R = 300 - 200 = 100 lbf Draw the SFDForces in the both the planes are at right angle, therefore, resultant at each location (from both the diagrams) can be writtNow, for AISI 1045 CD steel: Ultimate tensile strength, S = 91 ksi. Yield strength, S, = 77 ksi. Corrected endurance limit, SMean shear stress, T, 772 = .(a) ...... 167 ad 16x900 ad psi (b) 16T Alternating shear stress, ta = ad 16x900 ad psi SubstituFor combined stress problem from MSST, S Vo? + 47. y n 2 77 54.3056 ad 53.802 +4 ad 2.5 = 120530.98 ad d=1.075 in 77 2.5 (b)Similarly, 54.3056 ad ksi For combined stress problem from MSST, S yr +41 eg n 77 54.3056 ad +4 = 2.5 92.65 d (ď -0.48688) 3